Liquid is filled in a vessel which is kept in a room with temperature 20 C. When the temperature of the liquid is 80 C, then it looses heat at the rate of 60 cal/s What will be the rate of loss of heat when the temperature of the liquid is 40 C
 
  1. 180 cal/s
  2. 40 cal/s
  3. 30 cal/s
  4. 20 cal/s

Grade:12th pass

1 Answers

Eshan
5 years ago
Rate of loss of heat\proptodifference in temperature of body and the surrounding

\implies \dfrac{H_2}{H_1}=\dfrac{40-20}{80-20}=\dfrac{1}{3}

\implies H_2=20cal/s